please answe all questions carefully
pick any movie

Bell I Need To... Ochoose one of my favourite newly-released movies, and then find and record the box office revenues for the first 10 weeks of its run (this time period can be adjusted depending on the information that is available) Ograph the box office data from my research determine whether a single function can be used to describe the graph, or if it is m appropriate to use different functions for different parts of the domain Odevelop a function (or functions) to model the movie's cumulative box office revenue use my function to prediet the cumulative revenue after week 15 consider whether my model can be used to evaluate the box office revenue for all movies, and be prepared to present my findings to the class
PART 3: At the Movies . Investigate one of your favourite movies. Find and record the box office revenues for the first 10 weeks. You may wish to change the time period depending on the availability of data, but try to get about ten successive data points. Graph the data. Which type of function do you think would best describe the graph? Is one funetion appropriste or do you think it is more appropriate to use different functions for different parts of the dom? Develop a function (or functions) to model the movie's cumulative box office revenue. Use your function to predict the cumulative revenue after week 15 Discuss whether this model will work for all movies. Record a video of yourself giving a brief cresentation of your findings.
Answer #1

Inception is one of my favorite movies. Let's consider the box office collections for Inception for the first 10 weeks:

Week Weekly collections ($)
1 100158412
2 65670084
3 43303603
4 28052332
5 17008984
6 11450149
7 6929690
8 6836334
9 3753130
10 2643568

a) The graph for the weekly collections looks as follows:

350000000 300000000 y 8E+07n(x) + 1E+08.. 250000000 200000000 150000000 100000000 50000000 0 0 8 2 A 6 10 12 Week Cum. Collet

b) As seen above from the trendline fitted in Excel, an exponential function is appropriate to describe the graph.

c) As seen above, a function that can model the weekly collections is:

y = 8*10^7*ln(x) + 10^8

d) Cumulative revenue after Week 15 = 8*10^7*ln(15) + 10^8 = $ 316644016

e) This model might not work for all movies. For example, some of the movies become popular only after a few weeks (as it happened with Batman Begins). Hence, the cumulative collection curve might start really low and shoot up after a certain period, then dip suddenly. This is different from the collections modeled by the curve in this question.